# TSQCA: Threshold-Sweep Qualitative Comparative Analysis in R

### v2

@@ 1.3 Contributions of TSQCA
The package transforms thresholds from fixed parameters to analytical variables, enabling researchers to systematically map the solution space. This shift in perspective—from treating thresholds as “assumptions” to treating them as “objects of study”—opens new analytical possibilities.
@@ 2.2 The concept of hierarchical sufficiency
We introduce this concept as a descriptive pattern observed through threshold sweep analysis, not as a formal theoretical law.
@@ 2.3 Relationship to existing robustness research
Discussion of robustness in QCA has been developed most prominently by Skaaning (2011) and Oana and Schneider (2021). Related work on sensitivity diagnostics have sought to formalize how QCA results respond to parameter perturbations (Thiem et al., 2016). More recently, Baumgartner (2022) investigated the concept of robust sufficiency and, through simulation experiments, benchmarked the performance of different QCA solution types in recovering robustly sufficient conditions.
@@ 3.3 Relationship to the QCA package
Through this design, TSQCA functions as a “meta-analysis framework that repeatedly calls QCA.”
@@ 3.4 Package structure
|   |-- otSweep.R         # Outcome threshold sweep
|   |-- ctSweepS.R        # Single condition threshold sweep
|   |-- ctSweepM.R        # Multiple condition threshold sweep
|   |-- dtSweep.R         # Dual threshold sweep
|   |-- generate_report.R # Report generation
|   |-- utils.R           # Utility functions
|   ‘-- data.R            # Dataset documentation
@@ 3.6 Function overview
otSweep() Outcome threshold sweep sweep_range, thrX, include, dir.exp
ctSweepS() Single condition sweep sweep_var, sweep_range, thrY, dir.exp
ctSweepM() Multi-condition sweep sweep_list, thrY, include, dir.exp
dtSweep() Dual threshold sweep sweep_range_Y, sweep_list_X, dir.exp
@@ 4.1.2 Function signature
  include = "",
  dir.exp = NULL,
  return_details = FALSE
@@ 4.2.2 Function signature
  thrX_default,
@@
  include = "",
  dir.exp = NULL,
  return_details = FALSE
@@ 4.3.2 Function signature
  sweep_list,
@@
  include = "",
  dir.exp = NULL,
  return_details = FALSE
@@ 4.4.2 Function signature
  include = "",
  dir.exp = NULL,
  return_details = FALSE
@@ 7.2 Best practices
    Start with otSweep: Begin by exploring outcome threshold sensitivity before investigating condition thresholds.
@@
    Use meaningful threshold ranges: Select threshold ranges that correspond to substantively meaningful categories.
@@
    Document threshold rationale: Explicitly state why specific thresholds were chosen and what they represent substantively.
@@
    Report sensitivity results: Include threshold sensitivity analysis as a standard part of QCA reporting.
@@
    Interpret patterns, not just solutions: Focus on how solution structures evolve across thresholds, identifying recurring patterns such as hierarchical sufficiency.
@@ 8.1 Contributions
    Methodological innovation: We transform thresholds from fixed assumptions into objects of systematic study, enabling researchers to examine how causal structures emerge, disappear, and reorganize as thresholds vary.
@@ 8.2 Conceptual implications for QCA practice
The hierarchical sufficiency pattern illustrated in this paper highlights a descriptive relationship between outcome stringency and causal complexity. Higher outcome levels typically requires more demanding combinations of conditions, whereas broader explanatory coverage is associated with more modest targets. This relationship provides a framework for understanding how causal complexity varies with outcome stringency.
@@ 8.4 Concluding remarks
TSQCA represents a first step toward standardizing threshold sensitivity analysis in QCA research. By making such analyses accessible through user-friendly functions and structured outputs, we aim to elevate threshold examination from an occasional robustness check to a core component of QCA practice.
@@ 9 Computational details
R version 4.5.2 (2025-10-31)
Platform: aarch64-apple-darwin20
Running under: macOS Tahoe 26.2
@@
otSweep() Outcome threshold sweep dat, outcome, conditions, sweep_range, thrX, dir.exp
ctSweepS() Single condition sweep dat, outcome, conditions, thrY, sweep_var, sweep_range, dir.exp
ctSweepM() Multi-condition sweep dat, outcome, conditions, thrY, sweep_list, dir.exp
dtSweep() Dual threshold sweep dat, outcome, conditions, sweep_range_Y, sweep_list_X, dir.exp

### v3

@@ 1.3 Contributions of TSQCA
The package transforms thresholds from fixed parameters to analytical variables, enabling researchers to systematically map the solution space. This shift in perspective—from treating thresholds as “assumptions” to treating them as “objects of study”—opens additional analytical possibilities.
@@ 2.2 The concept of hierarchical sufficiency
We introduce this concept as a descriptive pattern observed through threshold sweep analysis, not as a formal theoretical law. The term does not imply a nested-set relationship in the strict set-theoretic sense, but rather describes an empirical pattern observed across threshold levels.
@@ 2.3 Relationship to existing robustness research
Discussion of robustness in QCA has been developed most prominently by Skaaning (2011) and Oana and Schneider (2021). Related work on sensitivity diagnostics has sought to formalize how QCA results respond to parameter perturbations (Thiem et al., 2016). More recently, Baumgartner (2022) investigated the concept of robust sufficiency and, through simulation experiments, benchmarked the performance of different QCA solution types in recovering robustly sufficient conditions.
@@ 3.3 Relationship to the QCA package
Through this design, TSQCA functions as a meta-analytic framework that repeatedly calls QCA.
@@ 3.4 Package structure
|   |-- TSQCA-package.R      # Package-level documentation
|   |-- tsqca_core.R         # Core utilities (binarization, validation)
|   |-- tsqca_ots_dts.R      # otSweep() and dtSweep()
|   |-- tsqca_cts.R          # ctSweepS() and ctSweepM()
|   |-- tsqca_report.R       # Report generation
|   |-- tsqca_config_chart.R # Configuration chart generation
|   |-- tsqca_helpers.R      # Helper functions
|   |-- tsqca_methods.R      # S3 print/summary methods
|   ‘-- sample_data.R        # Dataset documentation
@@ 3.6 Function overview
otSweep() Outcome threshold sweep sweep_range, thrX, pre_calibrated, include, dir.exp
ctSweepS() Single condition sweep sweep_var, sweep_range, thrY, pre_calibrated, dir.exp
ctSweepM() Multi-condition sweep sweep_list, thrY, pre_calibrated, include, dir.exp
dtSweep() Dual threshold sweep sweep_range_Y, sweep_list_X, pre_calibrated, dir.exp
@@ 4.1.2 Function signature
  pri.cut = 0,
  extract_mode = c("first", "all", "essential"),
  return_details = TRUE
@@ 4.1.3 Example with bundled sample data
4.1.4 Reading the output
The summary table contains one row per threshold value tested. The columns are interpreted as follows:
  • thrY: The outcome threshold value applied at that iteration.
  • expression: The minimized solution formula. The operator * denotes logical conjunction (AND), + denotes logical disjunction (OR), and a leading ~ denotes negation (absence of the condition). No solution indicates that no configuration met the consistency and frequency cutoffs at that threshold.
  • inclS: Solution consistency. Values of 0.8 or above are conventionally required for sufficiency claims (Schneider and Wagemann, 2012).
  • covS: Solution coverage. Higher values indicate greater empirical relevance of the solution.
  • n_solutions: The number of logically equivalent solutions returned by Boolean minimization. When this exceeds 1, the extract_mode argument controls how they are handled (see Section 5.2).
Three outcome patterns are commonly encountered:
Stable solution: The same expression appears across multiple threshold values, indicating that the finding is robust to threshold variation. In the example above, X3 + X1*X2 is returned at both thrY = 6 and thrY = 7.
Changing solution: Different expressions appear at different thresholds. This signals threshold sensitivity and warrants further investigation using ctSweepS() or ctSweepM() to identify the source of instability.
No solution: No configuration satisfies the cutoffs at the given threshold. This typically occurs when the threshold is set too high, leaving too few cases in the outcome set to support consistent configurations.
@@ 4.2.2 Function signature
  thrY,
  thrX_default = 7,
  pre_calibrated = NULL,
  dir.exp = NULL,
  include = "",
@@
  pri.cut = 0,
  extract_mode = c("first", "all", "essential"),
  return_details = TRUE
@@ 4.2.3 Example
4.2.4 Reading the output
The threshold column records the value applied to sweep_var at each iteration; all other condition thresholds remain fixed at thrX_default. The remaining columns (expression, inclS, covS, n_solutions) follow the same interpretation as described for otSweep() above.
In the example, the solution changes at threshold = 6: the term X3 disappears from the formula, leaving only X1*X2. This indicates that the threshold choice for X3 influences whether X3 enters the solution. At thresholds 7–9, the solution stabilizes, suggesting robustness across this range.
@@ 4.3.2 Function signature
  pre_calibrated = NULL,
  dir.exp = NULL,
  include = "",
@@
  pri.cut = 0,
  extract_mode = c("first", "all", "essential"),
  return_details = TRUE
@@ 4.3.3 Example
4.3.4 Reading the output
Each row corresponds to one threshold combination, identified by the threshold column (showing all condition threshold values) and a numeric combo_id. The solution columns are interpreted as before.
Scanning across rows, the analyst can identify which threshold combinations yield the same solution and which produce divergent results. In the example, combinations where X3 = 6 (combo_ids 1–4) tend to produce simpler solutions without X3, while combinations where X3 = 7 (combo_ids 5–8) more consistently include X3 in the solution. This pattern suggests that the X3 threshold is the primary driver of solution variation across the grid.
@@ 4.4.2 Function signature
  pri.cut = 0,
  extract_mode = c("first", "all", "essential"),
  return_details = TRUE
@@ 4.4.3 Example
4.4.4 Reading the output
The dtSweep() output extends the ctSweepM() layout with an additional thrY column, creating a two-dimensional view of the threshold space. Each unique combo_id corresponds to one configuration of condition thresholds; for each combo_id, multiple rows appear—one per outcome threshold value tested.
Comparing rows with the same combo_id across different thrY values reveals whether the solution is stable as the outcome threshold changes. Comparing rows with the same thrY across different combo_id values reveals condition-threshold sensitivity at a fixed outcome level. In the example, the solution X3 + X1*X2 at combo_id = 8 is consistent across both thrY = 6 and thrY = 7, indicating robustness in two dimensions simultaneously.
@@ 4.5 Notes on output structure and reproducibility
4.6 Mixed crisp/fuzzy analysis: the pre_calibrated argument (v1.3.0)
Version 1.3.0 introduces the pre_calibrated argument to all four sweep functions. This argument allows researchers to pass fuzzy-set membership scores—values in [0,1][0,1] produced by QCA::calibrate()—directly to QCA::truthTable() without binarization, while still sweeping other conditions on their original raw scale.
4.6.1 Motivation
In some research designs, theoretical considerations justify calibrating certain conditions as fuzzy sets (e.g., a “youth” set with a calibrated crossover point), while other conditions are more naturally treated on their original Likert or ratio scale within a threshold-sweep framework. Prior to v1.3.0, TSQCA required all conditions to be binarized via qca_bin(), which prevented this mixed workflow. The pre_calibrated argument removes this restriction without altering the behavior of existing code.
4.6.2 Usage
Variables listed in pre_calibrated are passed directly to QCA::truthTable() without binarization. All other conditions are binarized using qca_bin() at the thresholds specified in thrX, as usual. Pre-calibrated variables do not require a thrX entry, since no binarization is applied to them.
Sweep variables should be kept on their original raw scale. When a variable is swept on its raw scale (e.g., Likert scores 6, 7, 8), each threshold carries direct substantive meaning. Pre-calibrating a sweep variable into fuzzy membership values and then sweeping over those values (e.g., 0.3, 0.5, 0.7) makes interpretation indirect and harder to justify on substantive grounds.
4.6.3 Validation and warnings
The function raises an error if a variable listed in pre_calibrated is not found in conditions, or if its values fall outside [0,1][0,1] (which would indicate that a raw-scale variable was passed by mistake). A warning is issued if the variable contains NA values.
If a variable appears in both pre_calibrated and a sweep list (sweep_list_X in dtSweep(), or sweep_list in ctSweepM()), the function issues a warning and uses the pre-calibrated values, ignoring the sweep thresholds.
4.6.4 Backward compatibility
The default value pre_calibrated = NULL preserves v1.2.0 behavior exactly. All existing scripts run without modification.
@@ 7.2 Best practices
    Start with otSweep: Begin by exploring outcome threshold sensitivity before investigating condition thresholds. This establishes a baseline understanding of how the solution changes with the outcome definition and helps identify a stable outcome threshold range to fix for subsequent condition sweeps.
    # Step 1: single-value test to verify the setup
    result_test <- otSweep(
      dat = dat, outcome = "Y",
      conditions = c("X1", "X2", "X3"),
      sweep_range = 7,                  # single value first
      thrX = c(X1 = 7, X2 = 7, X3 = 7),
      include = "?", dir.exp = c(1, 1, 1)
    # Step 2: expand to the substantively meaningful range
    result_ots <- otSweep(
      dat = dat, outcome = "Y",
      conditions = c("X1", "X2", "X3"),
      sweep_range = 5:9,
      thrX = c(X1 = 7, X2 = 7, X3 = 7),
      include = "?", dir.exp = c(1, 1, 1)
@@
    Use meaningful threshold ranges: Select threshold ranges that correspond to substantively meaningful categories. For a 10-point Likert scale, the range 6:8 represents the upper portion of the scale; thresholds below 5 or above 9 are rarely theoretically defensible. Avoid sweeping the entire numeric range of a variable without substantive justification.
@@
[…]
    # Manageable: 2 x 2 x 2 = 8 combinations
    sweep_list_small <- list(X1 = 6:7, X2 = 6:7, X3 = 6:7)
    # Use with caution: 5 x 5 x 5 = 125 combinations
    sweep_list_large <- list(X1 = 5:9, X2 = 5:9, X3 = 5:9)
@@
    Document threshold rationale: Explicitly state why specific thresholds were chosen and what they represent substantively. A threshold of 7 on a 10-point satisfaction scale implies that cases scoring 7 or above are treated as “highly satisfied”; this should be justified on theoretical or measurement grounds rather than chosen to optimize fit.
@@
    Interpret patterns, not just solutions: Focus on how solution structures evolve across thresholds, identifying recurring patterns such as hierarchical sufficiency (Section 2.2). A solution that appears at every threshold is more credible than one that appears only at a single threshold. When solutions change across thresholds, the analyst should examine which conditions enter or exit the formula and whether this change is theoretically interpretable.
@@ 8.1 Contributions
    Methodological innovation: We reconceptualize thresholds from fixed assumptions as objects of systematic study, enabling researchers to examine how causal structures emerge, disappear, and reorganize as thresholds vary.
@@ 8.2 Conceptual implications for QCA practice
The hierarchical sufficiency pattern illustrated in this paper highlights a descriptive relationship between outcome stringency and causal complexity. Higher outcome levels typically require more demanding combinations of conditions, whereas broader explanatory coverage is associated with more modest targets. This relationship provides a framework for understanding how causal complexity varies with outcome stringency.
@@ 8.4 Concluding remarks
TSQCA represents a first step toward systematizing threshold sensitivity analysis in QCA research. By making such analyses accessible through user-friendly functions and structured outputs, we aim to elevate threshold examination from an occasional robustness check to a core component of QCA practice.
@@ 9 Computational details
R version 4.4.2 (2024-10-31 ucrt)
Platform: x86_64-w64-mingw32
Running under: Windows 11 x64
@@
otSweep() Outcome threshold sweep dat, outcome, conditions, sweep_range, thrX, pre_calibrated, dir.exp
ctSweepS() Single condition sweep dat, outcome, conditions, thrY, sweep_var, sweep_range, pre_calibrated, dir.exp
ctSweepM() Multi-condition sweep dat, outcome, conditions, thrY, sweep_list, pre_calibrated, dir.exp
dtSweep() Dual threshold sweep dat, outcome, conditions, sweep_range_Y, sweep_list_X, pre_calibrated, dir.exp